101 Questions for Pressure and Vapor Name Section

How do you solve these problems? Thank you
Answer #1

1)
it is asking for some experimental value. That you should be knowing after doing the experiment

2)
Hvap = 35.76 KJ/mol = 35760 J/mol

P1 = 62.53 mm
T1 = 22.5 oC = (22.5+273) K = 295.5 K
P2 = 597.4 mm
T2 = ?
use:
ln (P2/P1) = (Hvap/R)*(1/T1 - 1/T2)
ln(597.4/62.53) = (35760/8.314)*(1/295.5 - 1/T2)
(1/295.5 - 1/T2) = 5.247*10^-4
T2 = 350 K
= (350-273) oC
= 77 oC

Answer: 77 oC